Find all solutions for the equation Find+all+solutions+for+the+equation+3sec(4x)-3sqrt(2)=0
Answer by stanbon(75887)

3sec(4x)-3sqrt(2)=0
------
sec(4x) = sqrt(2)
-----
cos(4x) = sqrt(2)/2
-----
4x = +-(pi/4)
---
x = +-(pi/16)+-2npi
---------------------------
Cheers,
Stan H.
===================